# Unfitted finite element modelling of surface-bulk viscous flows in animal cells

### v1

@@ 3.1. Embedded geometry setup
Mesh and geometry configuration are illustrated in Figure 2(c). Let $\mathcal{T}_{h}$ be a quasi-uniform background grid in $\mathbb{R}^{3}$, composed of either $n$-simplices or $n$-cubes and with a characteristic mesh size $h=\max_{T\in\mathcal{T}_{h}}\{\mathrm{diam}(T)\}$. We assume that $\mathcal{T}_{h}$ meshes a bounding-box $\mathcal{B}$ of $\Omega$, covering but not necessarily fitting to the boundary $\Gamma$. To circumvent minor technical details, we require that $\mathcal{T}_{h}$ is fine enough to resolve the curvature of $\Gamma$. A precise statement of this assumption needs further formalism that we skip here, for the sake of conciseness; it can be usually found in specialised works such as [de2017condition, gurkan2019stabilized].
The sharp-interface representation is carried out with the level-set method [osher1988fronts]. Accordingly, the problem geometry is implicitly defined as the zero isosurface of a smooth level-set function $\phi(\boldsymbol{x}):\mathbb{R}^{3}\to\mathbb{R}$, that verifies $\Gamma=\{\boldsymbol{x}:\phi(\boldsymbol{x})=0\}$ and $\Omega=\{\boldsymbol{x}:\phi(\boldsymbol{x})<0\}$, see Figure 2(d). The level-set method is the simplest way to track surfaces without an explicit parametrisation. It is also a reasonable choice in our transient context, because surface dynamics are driven by the mechanics of the cortex –Equations (15)-(17). Therefore, in practice, we will only need to define the level set function $\phi^{0}$ of the initial cell configuration $\Gamma^{0}$, a simple sphere or spheroid, in most situations.
@@
Figure 3. Illustration of an unfitted FE approximation for a bulk PDE in $\Omega$. (a) The FE discretisation is built on the $\Omega$-active mesh $\mathcal{T}_{h}^{\Omega}$ formed by the grid cells intersecting $\Omega$. The simplest scenario is to approximate the problem with a linear Lagrangian FE space $\mathcal{V}_{h}^{\Omega}$, where the shape functions and degrees of freedom (in yellow circles) are associated to the vertices of $\mathcal{T}_{h}^{\Omega}$. (b) The FE solution $u_{h}$ of the problem is found in $\mathcal{T}_{h}^{\Omega}$ and its restriction $u_{h}{{}_{|}{{}_{\Omega}}}$ is the approximation to the continuous problem. (c) The small cut cell problem. Illustration of a badly cut cell $T$ with a very small cut portion in $\Omega$. If $a$ is a DOF of $\mathcal{V}_{h}^{\Omega}$, we see that its support inside $\Omega$ is so small that it can lead to ill-conditioning (see Equation (18)). (d) A way to fix the small cut cell problem, proposed in, e.g., the Aggregated FEM [badia2018aggregated], is to extrapolate the DOF value $u_{a}$ in terms of the DOF values of an interior cell $T(a)$. Purple arrows indicate the path mapping $a$ to $T(a)$ according to a cell aggregation scheme.
We readily see that discrete extensions provide a route to massively simplify the mesh generation step of FE problems in general, by enabling their resolution on a fixed grid. However, this comes at a high cost. Unfitted methods suffer from three main drawbacks: (1) numerical integration on cut cells requires dedicated procedures; (2) essential (e.g. Dirichlet) boundary conditions must be weakly imposed; and (3) naive discrete extensions often lead to (almost) singular linear systems, the so-called small cut-cell problem. All these issues and possible remedies have been extensively reviewed in, e.g., [de2023stability, schillinger2015finite].
Of all three challenges, the small cut-cell problem is widely viewed as the most critical one, involving the biggest effort for its mitigation or elimination [Burman2015, de2021numerical]. Let us give an intuitive, yet general idea of the issue. When working with moving boundaries and interfaces, it is often impractical, if not almost impossible, to tune with manual intervention how the grid overlaps the geometry. In particular, we must cope with any type of intersection between the grid cells and the geometry, including arbitrarily small ones, see example in Figure 3(c). In this situation, we may have basis functions with very reduced support inside the geometry. If using a classical FE method, these basis functions lead to rows and columns in the linear system with values close to zero, a clear case of ill-conditioning.
In fact, rigorous mathematical analysis for the bulk problem in a fixed-grid [de2017condition] proves that the condition number $\kappa$ of a standard FE system $\mathbf{A}$ for the Laplace operator satisfies
@@ 3.2.1. Bulk viscous flows
As there are no mesh nodes on the boundary $\Gamma$, it is not straightforward to enforce the Dirichlet condition of Equation (14) in the usual strong sense, that is, by prescribing the value of the condition at the boundary nodes. For this reason, we impose it in a weak sense, that is, by augmenting the variational formulaton with additional terms [fernandez2004imposing]. Usage of the widely employed Nitsche method [benzaken2024constructing, nitsche1971uber] gives rise to the forms
@@
In order to state the discrete problem, it remains to choose the FE space for the velocity-pressure pair $(\boldsymbol{u}_{h},p_{h})$. Our goal is to leverage the Lagrangian FE space $\mathcal{V}_{h}^{\Omega}$ in $\mathcal{T}_{h}^{\Omega}$ formed by the pair $\mathcal{Q}_{2}\times\mathcal{P}_{1}$ of (element-wise) continuous quadratic velocities and discontinuous linear pressures (restricted to zero mean global pressures). If we were in the body-fitted case, $\mathcal{V}_{h}^{\Omega}$ would be a classical choice, endowed with a discrete inf-sup stability condition [ern2004theory]. Unfortunately, $\mathcal{V}_{h}^{\Omega}$ does not preserve this key stability property in the unfitted case and, in presence of small cut elements, suffers from the ill-conditioning issues explained in Section 3.1.
A natural approach to avert the numerical instabilities is to eliminate from the linear system all DOFs linked to basis functions with too small support inside $\Omega$, subsequently referred to as ill-posed DOFs. This can be achieved, among others, with the Aggregated Finite Element Method (AgFEM[badia2018aggregated]. The main idea of AgFEM is to extrapolate the values of ill-posed DOFs in terms of the values of DOFs sitting on a cell with large-enough intersection with $\Omega$. For simplicity, we directly assume this cell is $\Omega$-interior, that is, strictly inside $\Omega$. The extrapolation amounts to enforce linear DOF constraints of the form
@@
We readily see that to define and evaluate the above constraints in $\mathcal{V}_{h}^{\Omega}$, we need to detect and gather all its ill-posed DOFs and assign to each one of them a suitable interior cell. For this, we resort to an automatic and easy-to-parallelise cell aggregation scheme [verdugo2019distributed], which derives this map from cell paths linking badly cut cells to interior cells. Note that, for good numerical accuracy and stability, cell paths should be as short as possible and connected through interior or cut facets [badia2018aggregated, Lemma 2.2].
Thanks to cell aggregation, we can apply Equation (19) to all ill-posed DOFs, pruning all basis functions associated with badly cut cells. This results in a restriction of the original FE space $\mathcal{V}_{h}^{\Omega}$ to the so-called aggregated FE space $\mathcal{V}_{h,\rm ag}^{\Omega}$. In this way, aggregated FE spaces fix the ill-conditioning issues of vanilla unfitted FEM; they are stable, regardless of the configuration of the cut cells, and retain the optimal approximation properties and condition number bounds of body-fitted FE methods. We refer to, e.g., [badia2018aggregated, burman2022cutfem] for fundamental theoretical analysis and results about AgFEM.
Our particular realisation of $\mathcal{V}_{h,\rm ag}^{\Omega}$ for the Stokes problem follows the robust and stable mixed formulation introduced in [Badia2018Mixed], opting for a serendipity extension and no pressure stabilisation. Hence, the discretisation of the bulk problem given by Equations (12)-(14) finally reads: Find $(\boldsymbol{u}_{h},p_{h})\in\mathcal{V}_{h,\rm ag}^{\Omega}$ such that
@@ 3.2.2. Surface viscous flows
with a very small friction term premultiplied by $\rho\!\ll\!1$ required, when $\Gamma$ is a closed surface, to eliminate Killing vector fields [sakai1996riemannian]. The right-hand side terms
@@
For surface problems, robustness to cut location is attained by perturbing the problem with a stabilisation term, in the spirit of trace FE methods [Olshanskii2017]. In particular, we adopt the normal derivative volume stabilisation, since it leads to optimal discretisation errors and condition number bounds for the vector Laplacian problem on fixed surfaces [jankuhn2021trace, burman2018cut]:
@@
where $\mathcal{N}_{h}^{\Gamma}=\bigcup_{T\in\mathcal{T}_{h}^{\Gamma}}\overline{T}$ is the domain covered by $\mathcal{T}_{h}^{\Gamma}$ and $\beta>0$ is a large-enough coefficient to ensure well-posedness. We observe that $S_{h}^{U}(\boldsymbol{U}_{h},\boldsymbol{V}_{h})$ implies integrating the basis functions in the whole cells cut by $\Gamma$, which necessarily eliminates the instabilities from small intersections of $\Gamma$ with $T\in\mathcal{T}_{h}^{\Gamma}$. On the other hand, the effect of this term is to extend the solution from $\Gamma$ to $\mathcal{N}_{h}^{\Gamma}$, by weakly imposing that the value on the surface is constant through the normal $\boldsymbol{n}_{\Gamma}$. In doing so, it crucially fixes another key problem of unfitted FEM for surface PDEs: the fact that bulk basis functions defined in $\mathcal{T}_{h}^{\Gamma}$ restricted on $\Gamma$ are, in general, linearly dependent. We refer to, e.g., [schollhammer2021higher, Figure 6] for a clear illustration of the linear dependency in bilinear quadrilateral cells.
@@ 3.2.3. Surface molecular transport
where $\gamma>0$ is a large-enough penalty coefficient. The particular choice of $S_{h}^{C}(C_{h}^{n},D_{h}^{n})$ is on the grounds of numerical stability and optimal error estimates proven in, e.g., [lehrenfeld2018stabilized]. Finally, the coupling term with $\boldsymbol{U}_{h}^{n}$ reads
@@
We observe that the above formulation conserves mass only approximately, due to the discretisation error in time and evaluating $C_{h}^{n-1}$ on $\Gamma^{n}$ (see end of Section 3.4). However, in the numerical experiments of Section 4, we notice that the error in mass conservation is reasonably controlled by the small time increments required to capture the problem nonlinearities. In other situations, we could require a conservative scheme as in, e.g. [olshanskii2014error, myrback2024high]. On the other hand, we use a SUPG variant [olshanskii2014stabilized] to stabilise high-activity cases ($\mbox{{Pe}}\gg 1$) that lead to advection-dominated transport.
@@ 3.2.4. Numerical integration
Except for the stabilisation terms, all integrals involved in the above weak formulations are defined in either $\Omega$ or $\Gamma$. Clearly, their evaluation on cut cells cannot be achieved with standard quadrature rules, because we need to restrict the integration to the portions $\Omega\cap T$ or $\Gamma\cap T$, $T\!\in\!\mathcal{T}_{h}^{\Gamma}$. There are many strategies to adapt the local integration on cut cells, reviewed in, e.g., [de2023stability, saye2022high]; they attempt, in most cases, to strike a good balance between accuracy and computational cost. In this work, we adopt the high-order quadrature rules introduced in [saye2022high] for curved surfaces and volumes implicitly defined by the level sets of multivariate polynomials. Underpinning our choice is the fact that shape dynamics are governed by the discrete FE unknown $\boldsymbol{U}_{h}$, which calls for the convenient representation of the evolving surface with piecewise polynomial level-sets (see next section).
@@ 3.3. Surface evolution
We outline now the evolving level-set method for shape evolution, as governed by Equations (15) and (17). First, we use the level-set description of the interface $\Gamma(\mathrm{t})$, introduced in Section 3.1, to recast the system into the classic initial value formulation [osher1988fronts]
@@
We observe that, at each time step, the rule above only involves updating the nodal values of the discrete level-set function. In order to evaluate $U_{\perp}^{\mathrm{e},n-1}$, we compute the value of $\boldsymbol{U}\cdot\boldsymbol{n}_{\Gamma}$ at the closest point projection to $\Gamma$ of each node of $\mathcal{V}_{h}^{\mathcal{B}}$. To this end, we resort to the high-order algorithms devised in [saye2014high]. In our numerical tests, we do not require to stabilise the scheme with, e.g., upwind methods, because the time steps are rather small. We further leverage closest point projections to represent $\phi^{n}$ as nodal projections of signed distance functions; in this way, we can assume $\|\nabla\phi^{n-1}\|\approx 1$.

### v2

@@ 3.1. Embedded geometry setup
Mesh and geometry configuration are illustrated in Figure 2(c). Let $\mathcal{T}_{h}$ be a quasi-uniform background grid in $\mathbb{R}^{3}$, composed of either $n$-simplices or $n$-cubes and with a characteristic mesh size $h=\max_{T\in\mathcal{T}_{h}}\{\mathrm{diam}(T)\}$. We assume that $\mathcal{T}_{h}$ meshes a bounding-box $\mathcal{B}$ of $\Omega$, covering but not necessarily fitting to the boundary $\Gamma$. To circumvent minor technical details, we require that $\mathcal{T}_{h}$ is fine enough to resolve the curvature of $\Gamma$. A precise statement of this assumption needs further formalism that we skip here, for the sake of conciseness; it can be usually found in specialised works such as [45, 66].
The sharp-interface representation is carried out with the level-set method [115]. Accordingly, the problem geometry is implicitly defined as the zero isosurface of a smooth level-set function $\phi(\boldsymbol{x}):\mathbb{R}^{3}\to\mathbb{R}$, that verifies $\Gamma=\{\boldsymbol{x}:\phi(\boldsymbol{x})=0\}$ and $\Omega=\{\boldsymbol{x}:\phi(\boldsymbol{x})<0\}$, see Figure 2(d). The level-set method is the simplest way to track surfaces without an explicit parametrisation. It is also a reasonable choice in our transient context, because surface dynamics are driven by the mechanics of the cortex –Equations (15)-(17). Therefore, in practice, we will only need to define the level set function $\phi^{0}$ of the initial cell configuration $\Gamma^{0}$, a simple sphere or spheroid, in most situations.
@@
Figure 3. Illustration of an unfitted FE approximation for a bulk PDE in $\Omega$. (a) The FE discretisation is built on the $\Omega$-active mesh $\mathcal{T}_{h}^{\Omega}$ formed by the grid cells intersecting $\Omega$. The simplest scenario is to approximate the problem with a linear Lagrangian FE space $\mathcal{V}_{h}^{\Omega}$, where the shape functions and degrees of freedom (in yellow circles) are associated to the vertices of $\mathcal{T}_{h}^{\Omega}$. (b) The FE solution $u_{h}$ of the problem is found in $\mathcal{T}_{h}^{\Omega}$ and its restriction $u_{h}{{}_{|}{{}_{\Omega}}}$ is the approximation to the continuous problem. (c) The small cut cell problem. Illustration of a badly cut cell $T$ with a very small cut portion in $\Omega$. If $a$ is a DOF of $\mathcal{V}_{h}^{\Omega}$, we see that its support inside $\Omega$ is so small that it can lead to ill-conditioning (see Equation (18)). (d) A way to fix the small cut cell problem, proposed in, e.g., the Aggregated FEM [9], is to extrapolate the DOF value $u_{a}$ in terms of the DOF values of an interior cell $T(a)$. Purple arrows indicate the path mapping $a$ to $T(a)$ according to a cell aggregation scheme described in Figure 4.
We readily see that discrete extensions provide a route to massively simplify the mesh generation step of FE problems in general, by enabling their resolution on a fixed grid. However, this comes at a high cost. Unfitted methods suffer from three main drawbacks: (1) numerical integration on cut cells requires dedicated procedures; (2) essential (e.g. Dirichlet) boundary conditions must be weakly imposed; and (3) naive discrete extensions often lead to (almost) singular linear systems, the so-called small cut-cell problem. All these issues and possible remedies have been extensively reviewed in, e.g., [44, 136].
Of all three challenges, the small cut-cell problem is widely viewed as the most critical one, involving the biggest effort for its mitigation or elimination [30, 43]. Let us give an intuitive, yet general idea of the issue. When working with moving boundaries and interfaces, it is often impractical, if not almost impossible, to tune with manual intervention how the grid overlaps the geometry. In particular, we must cope with any type of intersection between the grid cells and the geometry, including arbitrarily small ones, see example in Figure 3(c). In this situation, we may have basis functions with very reduced support inside the geometry. If using a classical FE method, these basis functions lead to rows and columns in the linear system with values close to zero, a clear case of ill-conditioning.
In fact, rigorous mathematical analysis for the bulk problem in a fixed-grid [45] proves that the condition number $\kappa$ of a standard FE system $\mathbf{A}$ for the Laplace operator satisfies
@@ 3.2.1. Bulk viscous flows
As there are no mesh nodes on the boundary $\Gamma$, it is not straightforward to enforce the Dirichlet condition of Equation (14) in the usual strong sense, that is, by prescribing the value of the condition at the boundary nodes. For this reason, we impose it in a weak sense, that is, by augmenting the variational formulaton with additional terms [54]. Usage of the widely employed Nitsche method [20, 109] gives rise to the forms
@@
In order to state the discrete problem, it remains to choose the FE space for the velocity-pressure pair $(\boldsymbol{u}_{h},p_{h})$. Our goal is to leverage the Lagrangian FE space $\mathcal{V}_{h}^{\Omega}$ in $\mathcal{T}_{h}^{\Omega}$ formed by the pair $\mathcal{Q}_{2}\times\mathcal{P}_{1}$ of (element-wise) continuous quadratic velocities and discontinuous linear pressures (restricted to zero mean global pressures). If we were in the body-fitted case, $\mathcal{V}_{h}^{\Omega}$ would be a classical choice, endowed with a discrete inf-sup stability condition [52]. Unfortunately, $\mathcal{V}_{h}^{\Omega}$ does not preserve this key stability property in the unfitted case and, in presence of small cut elements, suffers from the ill-conditioning issues explained in Section 3.1.
A natural approach to avert the numerical instabilities is to eliminate from the linear system all DOFs linked to basis functions with too small support inside $\Omega$, subsequently referred to as ill-posed DOFs. This can be achieved, among others, with the Aggregated Finite Element Method (AgFEM[9]. The main idea of AgFEM is to extrapolate the values of ill-posed DOFs in terms of the values of DOFs sitting on a cell with large-enough intersection with $\Omega$. For simplicity, we directly assume this cell is $\Omega$-interior, that is, strictly inside $\Omega$. The extrapolation amounts to enforce linear DOF constraints of the form
@@
We readily see that to define and evaluate the above constraints in $\mathcal{V}_{h}^{\Omega}$, we need to detect and gather all its ill-posed DOFs and assign to each one of them a suitable interior cell. For this, we resort to an automatic and easy-to-parallelise cell aggregation scheme [151], in Figure 4, which derives this map from cell paths linking badly cut cells to interior cells. Note that, for good numerical accuracy and stability, cell paths should be as short as possible and connected through interior or cut facets [9, Lemma 2.2].
Figure 4. Illustration of the cell aggregation scheme in $\mathcal{T}_{h}^{\Omega}$. (a) The goal of the scheme is to map every cut cell (light green) to a (unique) interior cell (dark green). This is achieved by iteratively growing cell aggregates (yellow borders) from interior to cut cells. (b) Given initial aggregates, corresponding to the interior cells, we find all facet-connected cut cell neighbours to the aggregates (white arrows). (c) Next, we merge all found cut cells into their aggregates and repeat the procedure until aggregating all cut cells into a single interior cell. (d) The cell paths (yellow arrows) formed in the aggregates yield the sought-after map. We refer to [9, 151] for the algorithmic description of the scheme.
Thanks to cell aggregation, we can apply Equation (19) to all ill-posed DOFs, pruning all basis functions associated with badly cut cells. This results in a restriction of the original FE space $\mathcal{V}_{h}^{\Omega}$ to the so-called aggregated FE space $\mathcal{V}_{h,\rm ag}^{\Omega}$. In this way, aggregated FE spaces fix the ill-conditioning issues of vanilla unfitted FEM; they are stable, regardless of the configuration of the cut cells, and retain the optimal approximation properties and condition number bounds of body-fitted FE methods. We refer to, e.g., [9, 33] for fundamental theoretical analysis and results about AgFEM.
Our particular realisation of $\mathcal{V}_{h,\rm ag}^{\Omega}$ for the Stokes problem follows the robust and stable mixed formulation introduced in [7], opting for a serendipity extension and no pressure stabilisation. Hence, the discretisation of the bulk problem given by Equations (12)-(14) finally reads: Find $(\boldsymbol{u}_{h},p_{h})\in\mathcal{V}_{h,\rm ag}^{\Omega}$ such that
@@ 3.2.2. Surface viscous flows
with a very small friction term premultiplied by $\rho\!\ll\!1$ required, when $\Gamma$ is a closed surface, to eliminate Killing vector fields [127]. The right-hand side terms
@@
For surface problems, robustness to cut location is attained by perturbing the problem with a stabilisation term, in the spirit of trace FE methods [113]. In particular, we adopt the normal derivative volume stabilisation, since it leads to optimal discretisation errors and condition number bounds for the vector Laplacian problem on fixed surfaces [81, 32]:
@@
where $\mathcal{N}_{h}^{\Gamma}=\bigcup_{T\in\mathcal{T}_{h}^{\Gamma}}\overline{T}$ is the domain covered by $\mathcal{T}_{h}^{\Gamma}$ and $\beta>0$ is a large-enough coefficient to ensure well-posedness. We observe that $S_{h}^{U}(\boldsymbol{U}_{h},\boldsymbol{V}_{h})$ implies integrating the basis functions in the whole cells cut by $\Gamma$, which necessarily eliminates the instabilities from small intersections of $\Gamma$ with $T\in\mathcal{T}_{h}^{\Gamma}$. On the other hand, the effect of this term is to extend the solution from $\Gamma$ to $\mathcal{N}_{h}^{\Gamma}$, by weakly imposing that the value on the surface is constant through the normal $\boldsymbol{n}_{\Gamma}$. In doing so, it crucially fixes another key problem of unfitted FEM for surface PDEs: the fact that bulk basis functions defined in $\mathcal{T}_{h}^{\Gamma}$ restricted on $\Gamma$ are, in general, linearly dependent. We refer to, e.g., [137, Figure 6] for a clear illustration of the linear dependency in bilinear quadrilateral cells.
@@ 3.2.3. Surface molecular transport
where $\gamma>0$ is a large-enough penalty coefficient. The particular choice of $S_{h}^{C}(C_{h}^{n},D_{h}^{n})$ is on the grounds of numerical stability and optimal error estimates proven in, e.g., [89]. Finally, the coupling term with $\boldsymbol{U}_{h}^{n}$ reads
@@
We observe that the above formulation conserves mass only approximately, due to the discretisation error in time and evaluating $C_{h}^{n-1}$ on $\Gamma^{n}$ (see end of Section 3.4). However, in the numerical experiments of Section 4, we notice that the error in mass conservation is reasonably controlled by the small time increments required to capture the problem nonlinearities. In other situations, we could require a conservative scheme as in, e.g. [112, 105]. On the other hand, we use a SUPG variant [111] to stabilise high-activity cases ($\mbox{{Pe}}\gg 1$) that lead to advection-dominated transport.
@@ 3.2.4. Numerical integration
Except for the stabilisation terms, all integrals involved in the above weak formulations are defined in either $\Omega$ or $\Gamma$. Clearly, their evaluation on cut cells cannot be achieved with standard quadrature rules, because we need to restrict the integration to the portions $\Omega\cap T$ or $\Gamma\cap T$, $T\!\in\!\mathcal{T}_{h}^{\Gamma}$. There are many strategies to adapt the local integration on cut cells, reviewed in, e.g., [44, 134]; they attempt, in most cases, to strike a good balance between accuracy and computational cost. In this work, we adopt the high-order quadrature rules introduced in [134] for curved surfaces and volumes implicitly defined by the level sets of multivariate polynomials. Underpinning our choice is the fact that shape dynamics are governed by the discrete FE unknown $\boldsymbol{U}_{h}$, which calls for the convenient representation of the evolving surface with piecewise polynomial level-sets (see next section).
@@ 3.3. Surface evolution
We outline now the evolving level-set method for shape evolution, as governed by Equations (15) and (17). First, we use the level-set description of the interface $\Gamma(\mathrm{t})$, introduced in Section 3.1, to recast the system into the classic initial value formulation [115]
@@
We observe that, at each time step, the rule above only involves updating the nodal values of the discrete level-set function. In order to evaluate $U_{\perp}^{\mathrm{e},n-1}$, we compute the value of $\boldsymbol{U}\cdot\boldsymbol{n}_{\Gamma}$ at the closest point projection to $\Gamma$ of each node of $\mathcal{V}_{h}^{\mathcal{B}}$. To this end, we resort to the high-order algorithms devised in [133]. In our numerical tests, we do not require to stabilise the scheme with, e.g., upwind methods, because the time steps are rather small. We further leverage closest point projections to represent $\phi^{n}$ as nodal projections of signed distance functions; in this way, we can assume $\|\nabla\phi^{n-1}\|\approx 1$.